AAVSOnet news, 2015-03-09

Mon, 03/09/2015 - 13:18

The telescopes that are operational include:  BSM_NM, BSM_Hamren, BSM_South, BSM_Berry, coker30, SRO, OC61.  They may have minor issues that have little impact on performance.

The telescopes that have problems at this time include: BSM_HQ (sticky roof) and TMO61 (communication issues between the telescope and computer).  We hope to have those problems resolved shortly.

A diffraction grating is getting installed on coker30, so that we will have three telescopes in the north with gratings: BSM_HQ, coker30 and TMLO61.  In the south, there are gratings on BSM_South and BSM_Berry.

Now that OC61 is functional again, we will resume testing of the eShel spectrograph.  OC61 has a two-position instument selector, so that both the direct imaging camera and the spectrograph can be mounted simultaneously.  They have different cameras, so currently a manual software change in MaximDL is needed to swap between modes.  This will probably mean that we will run either in queue-scheduled imaging mode or in spectroscopic mode for a given night, rather than switching between modes during the night.  A rapid switch is of course possible through manual intervention if a time-critical observation needs to be made.

As northern Spring rolls around, we will be bringing more options on-line.